find the domain of the function g(x)=10/4-5x
choose the correct domain
a. {x|x is a real number and x=10} ( # has a line thru it)
b. {x|x is a real number and > 4/5} the arrow has a line under it
c. {x|x is a real number and x # 4/5} ( # has a line thru it)
d. {x|x is a real number and x#0}( # has a line thru it)

Answer by ewatrrr(24785)


Hi
finding the domain of the function
g(x)=10/(4-5x) |Note use of ()
denominator CANNOT = 0 IF 4-5x = 0 ... then x = 4/5
c. {x|x is a real number and x ≠ 4/5}